A uniform electric field exists in the region between two oppositely charged plane parallel plates. A proton is released from rest at the surface of the positively charged plate and strikes the surface of the opposite plate, 1.80 cm distant from the first, in a time interval of 2.30×10−6 s
Answer #1

Solution,

A)

Using 2nd equation of motion,

S = ut + 0.5 at^2

0.018 = 0 + 0.5 x a x (2.3 x 10^-6)^2

Acceleration, a = 6.81 x 10^9 m/s^2

Force, F = 1.67 x 10^-27 x 6.81 x 10^9 = 1.14 x 10^-17 N

Electric field, E = (1.14 x 10^-17)/(1.6 x 10^-19) = 71.03 N/C

B)

V = u + at = 0 + (6.81 x 10^9)(2.3 x 10^-6) = 15663 m/s

Plz do rate my answer.....